OpenStudy (anonymous):

HELP PLEASE The height of a tower is three times the height of a certain building. If the tower is 50 m taller than the building, how tall is the tower?

OpenStudy (kkutie7):

t=tower height b=building height t=3*b t=50m+b a way to think of it.

OpenStudy (anonymous):

OKAYYY so what do i do?

OpenStudy (kkutie7):

b=t/3 t=50+t/3 t-50=t/3 3t-150=t -150=t-3t -150=-2t 150=2t 75m=t I think this is right. you might want to check it out.

OpenStudy (anonymous):

this is really confusing

OpenStudy (kkutie7):

ok let me break it down for you

OpenStudy (anonymous):

alright

OpenStudy (kkutie7):

t=tower height b=building height t=3*b because The height of a tower is three times the height of a certain building t=50m+b because tower is 50 m taller than the building are you with me so far?

OpenStudy (anonymous):

yep so far so good!

OpenStudy (kkutie7):

alright good =) what I did was take one of the equations ad solve for b. \[t=3*b\] \[\frac{t}{3}=b\]

OpenStudy (anonymous):

aye wait!

OpenStudy (kkutie7):

yes?

OpenStudy (anonymous):

wouldnt it be 3b=t

OpenStudy (kkutie7):

no because we are solving for b not t

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

but i dont understand how you got t over 3?

OpenStudy (kkutie7):

ok so t=3xb right? so in order to get b on one side of the equal sign by itself you need to divide both sides by 3 so \[\frac{t}{3}=\frac{3b}{3}\]

OpenStudy (anonymous):

whered you get the (x) after the 3?

OpenStudy (kkutie7):

that's the same as * just multiplication =)